Determine the measure of the vertical angles indicated. 2x + 46 and 3x + 10.
Answer by RAY100(1637)

Remember vertical angles are equal
.
2x+46 = 3x +10
.
36 = x
.
.
check
2(36) +46 = 3(36) +10
.
72 +46 = 108 +10
.
118 = 118,,,,,ok
.